Fermi National Laboratory


Accelerator Update

Wednesday March 3
The day shift began with Operations monitoring store 3271, with MiniBooNE experts conducting studies, and with the Recycler stashing antiprotons.

At 8:54 AM, Recycler shots stopped and the Antiproton Source (Pbar) returned to stacking.

Booster experts began a thirty-minute cogging study at 1:21 PM.

The Main Injector (MI) suffered a vacuum burst at 2:18 PM. Everything reset okay.

A Booster expert began tuning on the low-level RF system at 2:48 PM.

At 3:39 PM, Operations began sending beam to the experiment at Meson Center (MCenter).

A Meson power supply (MC1D) developed a problem 10:24 PM. Its output current dropped. Operators managed to bring the supply back up to its proper current, but it will need work later.

Thursday March 4
The midnight shift began with Operations monitoring store 3271, with Pbar stacking, with MiniBooNE taking beam, and with a MI expert conducting studies.

At 12:01 AM, Operations turned off beam to MCenter.

The MI studies finished at 1:08 AM.

Operations terminated store 3271 at 6:04 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 6:36 AM.

At 9:31 AM, Operations established store 3273 with an initial luminosity of 47.82E30.

Operators had tried to resume stacking, but noticed, at 9:38 AM, that the Debuncher wasn't getting any RF. Operators contacted an expert who, within ten minutes, had found and fixed a bad connector.

Power supply technicians replaced a bad reference module on MC1D at 10:36 AM.

At 12:44 PM, Operations began sending beam to Switchyard (SY).

By 4 PM, experts in MI, Recycler, SY, and MiniBooNE were conducting studies.

Operations began sending beam to MCenter at 4:36 PM. They ran until midnight.

Friday March 5
The midnight shift began with Operations monitoring store 3273, with Pbar stacking, with MiniBooNE taking beam, and with MI conducting studies.

A Booster RF station (BRF1) tripped off at 12:38 AM and then again three hours later. It successfully reset both times.

A Linac RF station (LRF3) tripped off twice; first at 2:46 AM and then about four hours later. It reset both times okay.

MI studies finished at 2:47 AM.

Today's Plans
The plans for today are to end the store around noon and then let the D-zero experimenters access their collision hall to search their detector for an electrical noise problem. They guess that this access may take anywhere from three to twelve hours. The machines, other than the TeV, will conduct studies during this time. Hopefully, this weekend will consist of stacking and stores.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/5/2004   email Fermilab